# HIRE: A Hybrid Learned Index for Robust and Efficient Performance under Mixed Workloads

### v1

@@ 5.5. Performance on Scalability
Figure 14. Effectiveness of Hybrid Nodes
Figure 15. Effectiveness of Non-Blocking Recalibration
@@ 5.7.2. Effectiveness of Non-Blocking Recalibration
To evaluate our non-blocking, cost-driven recalibration design, we study a modified HIRE variant which replaces our proposed mechanism with a simplified, single-threaded, and blocking recalibration process. The results on the OSM dataset, shown in Figure 15, indicate that without our non-blocking recalibration design, the HIRE variant exhibits significant latency spikes. The tail latency at the 99th and 99.9th percentiles increases by $2.9\times$ and $4.2\times$, respectively, compared to the full HIRE. These spikes result from large-scale structural adjustments, such as when a large model-based leaf node splits into multiple leaf nodes, forcing the index to stall for an extended period for structure modifications. This experiment confirms that our novel non-blocking recalibration mechanism is crucial for mitigating tail latency and ensuring robust performance, particularly under dynamic workloads.
@@
5.8. Performance on Concurrency
Although the initial design of HIRE did not support fully concurrent read-write operations, we observe that its structural modifications are localized, triggered only when legacy leaf nodes undergo splits, merges, or redistributions. This property allows us to develop a concurrent-safe version, denoted HIRE+, by incorporating a concurrency control mechanism similar to that of ALEX+ and LIPP+ in GRE (wongkham2022updatable). This approach augmented the existing RCU mechanism in HIRE with node-level locking to ensure thread-safe access. We evaluate the performance of HIRE+ under concurrent write-heavy workloads against other concurrent indexes, including B+-tree-OLC, ALEX+, and LIPP+. As PGM lacks a concurrent implementation, we replace it with FINEdex (li2021finedex), which is natively designed for concurrency. All baseline implementations are sourced from the GRE framework (wongkham2022updatable).
@@ 6. Related Work
The concept of the learned index, RMI (kraska2018case), emerged in 2018 as a static approach. Early research on learned indexes mainly focuses on static workload. Several works, such as RadixSpline (kipf2020radixspline), PLEX (stoian2021towards), and Shift-Table (hadian2021shift), demonstrate the potential of model-based indexing. However, a common criticism of these early designs is their inability to effectively support dynamic updates.
@@
To overcome the limitation, updatable learned indexes have been introduced, including FITing-Tree (galakatos2019fiting), PGM, and ALEX (ding2020alex). Both FITing-Tree and PGM employ a balanced structure built bottom-up. ALEX, another notable updatable learned index, extends the RMI concept by constructing the index top-down. It sacrifices balance and a bounded structure to achieve better model fitting at different data granularities, and linearizes the data by adding gaps into segments to create a sparser structure. These pioneering updatable learned indexes delineate a design space characterized by one machine learning model (linear model), two construction methods (top-down and bottom-up), and three buffer types (index-level, segment-level, and data-level). Along with benchmarks like SOSD (kipf2019sosd) for performance evaluation on real-world data, these approaches have triggered a “Cambrian explosion” of learned index proposals.
@@
During this surge, numerous learned indexes have been developed, revealing three main design trends. First, some works focus on optimizing model accuracy and structure for faster lookups. For example, LIPP achieves near error-free performance using a model-based insertion and splitting approach, although it sacrifices data locality. DILI (li2023dili) employs a concise, computation-efficient linear regression model for improved efficiency, while Hyper (zhang2024hyper) combines bottom-up and top-down construction approaches to balance latency and memory usage. SALI (ge2023sali) maintains internal statistical information to further refine its model. Second, other designs leverage modern hardware features to boost performance. For instance, Xindex (tang2020xindex) and FINEdex (li2021finedex) are designed to support concurrent operations. CARMI (zhang2021carmi) is the first learned index to fully exploit CPU cache characteristics. (zhong2022learned) extends optimizations to GPUs. Third, a shift toward case-specific learned indexes has emerged to address the limitations of general-purpose designs. Most learned indexes, including our proposed HIRE and others like ALEX, PGM and LIPP, are designed for numeric data, as modeling the CDF of strings for key lookup is a non-trivial challenge. This has led to the development of specialized solutions. For string data, SIndex (wang2020sindex) and LITS (yang2024lits) represent notable recent progress. For other use cases, indexes like FLIRT (yang2023flirt) and SWIX (liang2024swix) have been designed specifically for indexing streaming windows.
@@
There are also research efforts in optimizing learned indexes, including improving persistent memory efficiency (lu2021apex; cui2023learned; wang2024wipe; lan2024fully; lan2023simple; zhang2024making) and parameter tuning for optimal performance (chockchowwat2022tuning; chockchowwat2023airindex; wang2025new).

### v2

@@ 5.2.1. Point Lookup Query
Figure 9. Throughput on Range Queries
Figure 10. Tail Latency on Different Datasets
@@ 5.4. Performance on Different Match Rates
Figure 12. Throughput and Index Size on Different Datasizes
Figure 13. Index Build Time on Different Datasizes
@@ 5.7.2. Effectiveness of Non-Blocking Recalibration
To evaluate our non-blocking, cost-driven recalibration design, we study a modified HIRE variant which replaces our proposed mechanism with a simplified, single-threaded, and blocking recalibration process. The results on the OSM dataset, shown in Figure 15, indicate that without our non-blocking recalibration design, the HIRE variant exhibits significant latency spikes. Tail latency at the 99th and 99.9th percentiles increases by $2.9\times$ and $4.2\times$, respectively, compared to the full HIRE. These spikes result from large-scale structural adjustments, such as when a large model-based leaf node splits into multiple leaf nodes, forcing the index to stall for an extended period for structure modifications. This experiment confirms that our novel non-blocking recalibration mechanism is crucial for mitigating tail latency and ensuring robust performance, particularly under dynamic workloads.
Figure 14. Impact of Hybrid Nodes
Figure 15. Impact of Non-blocking Recalibration
5.8. Performance on Concurrency
Although the initial design of HIRE did not support fully concurrent read-write operations, we observe that its structural modifications are localized, triggered only when legacy leaf nodes undergo splits, merges, or redistributions. This property allows us to develop a concurrent-safe version, denoted HIRE+, by incorporating a concurrency control mechanism similar to that of ALEX+ and LIPP+ in GRE (Wongkham et al., 2022). This approach augmented the existing RCU mechanism in HIRE with node-level locking to ensure thread-safe access. We evaluate the performance of HIRE+ under concurrent write-heavy workloads against other concurrent indexes, including B+-tree-OLC, ALEX+, and LIPP+. As PGM lacks a concurrent implementation, we replace it with FINEdex (Li et al., 2021), which is natively designed for concurrency. All baseline implementations are sourced from the GRE framework (Wongkham et al., 2022).
@@ 6. Related Work
The concept of the learned index, RMI (Kraska et al., 2018), emerged in 2018 as a static approach. Early research on learned indexes mainly focuses on static workload. Several works, such as RadixSpline (Kipf et al., 2020), PLEX (Stoian et al., 2021), and Shift-Table (Hadian and Heinis, 2021), demonstrate the potential of model-based indexing. However, a common criticism of these early designs is their inability to effectively support dynamic updates.
@@
To overcome the limitation, updatable learned indexes have been introduced, including FITing-Tree (Galakatos et al., 2019), PGM, and ALEX (Ding et al., 2020). Both FITing-Tree and PGM employ a balanced structure built bottom-up. ALEX, another notable updatable learned index, extends the RMI concept by constructing the index top-down. It sacrifices balance and a bounded structure to achieve better model fitting at different data granularities, and linearizes the data by adding gaps into segments to create a sparser structure. These pioneering updatable learned indexes delineate a design space characterized by one machine learning model (linear model), two construction methods (top-down and bottom-up), and three buffer types (index-level, segment-level, and data-level). Along with benchmarks like SOSD (Marcus et al., 2020) for performance evaluation on real-world data, these approaches have triggered a “Cambrian explosion” of learned index proposals.
@@
During this surge, numerous learned indexes have been developed, revealing three main design trends. First, some works focus on optimizing model accuracy and structure for faster lookups. For example, LIPP achieves near error-free performance using a model-based insertion and splitting approach, although it sacrifices data locality. DILI (Li et al., 2023) employs a concise, computation-efficient linear regression model for improved efficiency, while Hyper (Zhang et al., 2024c) combines bottom-up and top-down construction approaches to balance latency and memory usage. SALI (Ge et al., 2023) maintains internal statistical information to further refine its model. Second, other designs leverage modern hardware features to boost performance. For instance, Xindex (Tang et al., 2020) and FINEdex (Li et al., 2021) are designed to support concurrent operations. CARMI (Zhang and Gao, 2022) is the first learned index to fully exploit CPU cache characteristics. (Zhong et al., 2022) extends optimizations to GPUs. Third, a shift toward case-specific learned indexes has emerged to address the limitations of general-purpose designs. Most learned indexes, including our proposed HIRE and others like ALEX, PGM and LIPP, are designed for numeric data, as modeling the CDF of strings for key lookup is a non-trivial challenge. This has led to the development of specialized solutions. For string data, SIndex (Wang et al., 2020) and LITS (Yang and Chen, 2024) represent notable recent progress. For other use cases, indexes like FLIRT (Yang et al., 2023) and SWIX (Liang et al., 2024) have been designed specifically for indexing streaming windows.
@@
There are also research efforts in optimizing learned indexes, including improving persistent memory efficiency (Lu et al., 2021; Cui et al., 2024; Wang et al., 2024; Lan et al., 2024, 2023; Zhang et al., 2024a) and parameter tuning for optimal performance (Chockchowwat, 2022; Chockchowwat et al., 2023; Wang et al., 2025).
@@ 7. Conclusion
Acknowledgements.
We are deeply thankful to the anonymous reviewers for their constructive reviews. This work is supported by Hong Kong RGC Grants (Project No. 12202024 and 12200022), the Swiss State Secretariat for Education, Research and Innovation (SERI) for the Prodasys project, and a European Research Council (ERC) Advanced Grant. Jianliang Xu is the corresponding author.